OpenStudy (fairy733189):

The expression 600(1.004274)4x shows the quarterly interest rate Markey earns after x years. Which expression shows the x isolated in the exponent in an equivalent form? 600(4.017096)x 600(0.251069)x 600(1.001067)x 600(1.017206)x

OpenStudy (fairy733189):

All of the x's are exponents

OpenStudy (fairy733189):

@vishweshshrimali5

OpenStudy (fairy733189):

Also the 4x is an exponent

OpenStudy (vishweshshrimali5):

Okay.. so \(600(1.004274)^{4x}\)

OpenStudy (fairy733189):

yes

OpenStudy (vishweshshrimali5):

Okay.. recall the exponents law... \(x^{mn} = (x^m)^n\)

OpenStudy (vishweshshrimali5):

So you can write 4x = 4 * x... any ideas what we can do next?

OpenStudy (fairy733189):

Im not really sure I don't think I remember learning this yet

OpenStudy (vishweshshrimali5):

We learn something new every day ;) ... So, let me give you another example...

OpenStudy (vishweshshrimali5):

Suppose I want to write...\(2^{3x}\) in terms of exponent x only

OpenStudy (fairy733189):

ok

OpenStudy (vishweshshrimali5):

Now, I apply the exponents law... \(2^{3x} = 2^{3*x} = (2^3)^x\) Clear so far?

OpenStudy (fairy733189):

would it be (2^3) ^x

OpenStudy (vishweshshrimali5):

Yeps!! Perfect

OpenStudy (vishweshshrimali5):

So, now, what is 2^3

OpenStudy (fairy733189):

great

OpenStudy (fairy733189):

8

OpenStudy (vishweshshrimali5):

Perfect! So we get \(8^x\)

OpenStudy (fairy733189):

ok

OpenStudy (vishweshshrimali5):

Now, can you solve your problem ?

OpenStudy (fairy733189):

let me try it will probably take about 5 minutes

OpenStudy (vishweshshrimali5):

No problem... take your time

OpenStudy (fairy733189):

would it be the first one?

OpenStudy (mathmale):

No. You've multiplied. That 4 is an exponent, not a multiplicand.

OpenStudy (fairy733189):

oh so it would be the last one

OpenStudy (fairy733189):

because you multiply the number in the parenthesis by itself 4 times

OpenStudy (vishweshshrimali5):

Perfect!!

OpenStudy (fairy733189):

Thank you soooo much. I am not that great at math but your helping me understand it a lot better.

OpenStudy (fairy733189):

I just have one more problem that I don't understand do you think you can help me with this last one?

OpenStudy (vishweshshrimali5):

:) Glad to be of help